# Evaluating the Effect of Retrieval Augmentation on Social Biases

### v1

@@ Abstract
Retrieval Augmented Generation (RAG) has gained popularity as a method for conveniently incorporating novel facts that were not seen during the pre-training stage in Large Language Model (LLM)-based Natural Language Generation (NLG) systems. However, LLMs are known to encode significant levels of unfair social biases. The modulation of these biases by RAG in NLG systems is not well understood. In this paper, we systematically study the relationship between the different components of a RAG system and the social biases presented in the text generated across three languages (i.e. English, Japanese and Chinese) and four social bias types (i.e. gender, race, age and religion). Specifically, using the Bias Question Answering (BBQ) benchmark datasets, we evaluate the social biases in RAG responses from document collections with varying levels of stereotypical biases, employing multiple LLMs used as generators. We find that the biases in document collections are often amplified in the generated responses, even when the generating LLM exhibits a low-level of bias. Our findings raise concerns about the use of RAG as a technique for injecting novel facts into NLG systems and call for careful evaluation of potential social biases in RAG applications before their real-world deployment.
@@
LLMs are trained on vast collections of texts typically sourced from the internet. These models encapsulate a broad spectrum of information, yet they fail to incorporate emerging facts after pre-training, leading to inaccuracies and hallucinatory outputs (Song et al., 2024; Niu et al., 2024; Agrawal et al., 2024). Traditional approaches to update LLMs with new information include continual pre-training Ke et al. (2022) and supervised fine-tuning Ouyang et al. (2022). However, updating parameters of LLMs with large datasets is time consuming and expensive even with parameter efficient methods Hu et al. (2021). Moreover, closed models such as GPT-4o restrict model parameter access.
RAG Lewis et al. (2020); Edge et al. (2024) offers a popular alternative by integrating real-time retrieval of documents to supplement the training data Izacard and Grave (2021); Jiang et al. (2024); Shuster et al. (2021). This approach allows LLMs to access and utilise information unavailable during their initial training.
@@
The document sets used in RAG are crucial as they directly influence the generated content. The inherent social biases of these documents, coupled with those encoded by the LLMs, determine the bias level of the outputs. Despite extensive evaluation of RAG systems for retrieval efficacy Wu et al. (2024); Laban et al. (2024); Yang et al. (2024) and factual accuracy Krishna et al. (2024); Soman and Roychowdhury (2024), their role in propagating social biases has been under explored.
This paper addresses this oversight by investigating how RAG influences social biases when LLMs are presented with externally sourced contexts, potentially laden with stereotypes. We analyse the bias propagation using BBQ Parrish et al. (2022), a QA-structured benchmark that assesses social biases in LLMs, across gender, age, race and religion applying three retrieval methods. Furthermore, we extend our analysis to include multilingual social bias evaluations in English, Japanese and Chinese. Our findings highlight several key points:
@@
    We find that all types of social biases are amplified when stereotypically biased documents are used for RAG. This is particularly worrying because despite the careful bias mitigations conducted prior to releasing LLMs, RAG can easily generate socially biased responses from those LLMs. Interestingly, we see that the level of social bias increment in larger LLMs tend to be smaller compared to that in smaller LLMs.
@@
    Overall, we find that social biases are affected to a lesser degree by the retrieval method used in RAG, while sparse retrieval methods tend to be more sensitive to social biases than dense retrieval methods. Surprisingly, social biases do not necessarily increase with the number of documents retrieved, demonstrating a trade-off due to the decreasing relevance of the documents to the input query.
@@
    This bias amplification is not confined to English and can be observed for non-English languages as well such as Japanese and Chinese, demonstrating a global challenge across languages.
We advocate for a reconsideration of how social biases are evaluated in RAG systems.
@@
Social Biases in LLMs:
LLMs are typically trained on extensive text collections sourced from the internet, which often contains various types of social biases that are then mirrored in the models’ behaviour Penedo et al. (2024). These biases can be assessed through two primary methods: intrinsic and extrinsic Goldfarb-Tarrant et al. (2021); Cao et al. (2022) evaluation. Intrinsic measures focus on biases within word embeddings or model predictions Caliskan et al. (2017); Nangia et al. (2020); Nadeem et al. (2021a); Kaneko et al. (2022a), while extrinsic measures analyse biases in outputs from downstream tasks such as Natural Language Inference (NLI) or question answering Webster et al. (2020); De-Arteaga et al. (2019).
@@
Although social biases of LLMs have been studied extensively for various downstream applications, the effect of RAG on NLG has been less frequently explored. Hu et al. (2021) proposed a three-level threat model and studied the sensitivity of RAG to the external datasets used for the retrieval. They found that the fairness of a RAG system can get easily compromised due to the social biases in the datasets used. Unlike their approach which uses a limited set of contexts from the BBQ dataset, our work incorporates a broader range of documents from diverse datasets, enhancing the generalisability of our findings.
Wu et al. (2025) explored fairness within RAG systems by examining disparities in retrieval performance between protected and non-protected groups, using data from FairRanking Track (Ekstrand et al., 2023) that focuses on protected attributes like binary gender (i.e. female vs. males) and geographic origin (i.e. non-Europeans vs. Europeans). This study primarily addressed fairness, defined as equitable retrieval performance, whereas our study extends the evaluation to social biases in multilingual contexts, providing a deeper understanding of bias dynamics in RAG beyond just fairness.
Figure 2: Overview of our RAG social bias evaluation protocol. Given a collection of documents, each document is encoded using an external encoder $f$ and a vector index is created over the collection of the documents. We use a question, paired with its ambiguous or disambiguated context, selected from the BBQ dataset as the query for the retrieval system. We then retrieve the top $k$ nearest neighbour documents to the query from the vector index, and provide them to the generator LLM, $g$, alongside with the question and the context. The generator is instructed to select the most suitable answer from given choices.
@@
§ 3.1 Background – RAG
Before we describe our social bias evaluation protocol for RAG, let us briefly describe the main components of a typical RAG system and how social biases could potentially influence each component. A typical RAG system is shown in Figure 2 and consists of several components.
Document Collection $\mathcal{D}$:
A RAG system is given an external document collection, $\mathcal{D}$, containing information (possibly was not available) in the pre-trained LLMs.
Retriever:
Documents are indexed to efficiently retrieve those most relevant to a given query. Both sparse and dense retrieval methods can be used for this purpose. In sparse retrieval, each document is tokenised and represented as a bag-of-tokens. Next, an inverted index is created over the set of documents. Similarly, the query is tokenised and under the conjunctive matching, documents that contain all of the tokens in the query are retrieved from the index. Dense retrieval employs a pre-trained text encoder Xu et al. (2023); Gao et al. (2021), $f$ to represent the query as well as each of the documents in $\mathcal{D}$ in a fixed $m$-dimensional embedding $f(d)(\in\mathbb{R}^{m})$ space. Next, a vector index is created over the embedded documents and approximate nearest neighbour (ANN) retrieval methods Malkov and Yashunin (2020); Guo et al. (2020) are used to efficiently retrieve similar documents to the query.
Generator $g$:
For a given query $q$, an LLM is used as the generator $g$ to produce a response, $\hat{r}(q)=g(q,d^{(q)}_{1},d^{(q)}_{2},\ldots,d^{(q)}_{k},I)$ using the set of top-$k$ retrieved documents for $q$, $\{d^{(q)}_{1},d^{(q)}_{2},\ldots,d^{(q)}_{k}\}$, and an instruction, $I$, that dictates the response formulation.
@@
To describe our social bias evaluation protocol for RAG, let us consider the example shown in Figure 2, selected from the BBQ dataset, where given the ambiguous context “We couldn’t start the board meeting at 9 am today because a man and a woman were late”, we must answer the question, “Who was the secretary?”. The query here is the concatenation of the question and the context, and relevant documents will be retrieved from the document collection. Next, the retrieved documents, question and the context will be shown to the generator, instructing with a prompt that lists three answer choices: (a) the man is the secretary, (b) the woman is the secretary, and (c) unknown. The generator is specifically instructed to select one of the three answer choices and return the letter corresponding to the selected answer.
@@
An LLM that is stereotypically biased towards the male gender would incorrectly predict (b), assuming that all secretaries are females. The choice that would not reflect an unfair gender bias in this example would be (c). However, note that an LLM which always selects (c) and refrain from answering any socially sensitive questions would appear to be unbiased, when in fact it might not be the case. Therefore, the BBQ dataset includes disambiguated contexts, which explicitly provide information to decide the correct answer without having to rely on any social biases encoded in the generator LLM or the retrieved documents. In the previous case, the disambiguation context provided in BBQ is “We had to wait for the woman because they were running the meeting, and the man was responsible for taking all the notes”. Given this disambiguated context the correct answer to this question would be (a).
Evaluating social biases under a RAG setting is particularly challenging for two main reasons.
  1. 1.
    Component Interaction: Each component (document collection, retriever and generator) in a RAG system can independently and collectively influence social bias propagation. In order to conduct a systematic and reproducible evaluation without conflating multiple factors, it is important to vary only one of the components, while keeping the others fixed.
  2. 2.
    Open-ended Generation: Automatically evaluating social biases under an open-ended generation setting is difficult because the same social bias can be expressed in different ways in the generator responses Esiobu et al. (2023). Modelling social bias evaluation in RAG as a multiple choice question-answering task enables us to evaluate social biases without having to consider open-ended generations.
Next, we discuss how social biases can influence each of the RAG components.
Biases in the Documents:
If there are many documents that express various levels of stereotypical social biases, then a subset of those documents can be retrieved even when the query does not explicitly mention any social biases. Revisiting our previous example, if there are many documents that mention females as secretaries in the document collection, it is possible that we will retrieve some of those biased documents, which could in return influence the generator to produce a biased response. We evaluate the effect of four types of social biases (i.e. gender, age, race, religion) (§ 4.2) in the document collection using three benchmark datasets covering English, Japanese Yanaka et al. (2024) and Chinese Huang and Xiong (2024) languages (§ 4.4). Moreover, as control settings we consider document collections that consists purely of stereotypical or anti-stereotypical documents in § 4.2.
Biases in the Retriever:
The text encoders used for embedding documents and queries for dense retrieval can also encode unfair social biases Bolukbasi et al. (2016); Kaneko et al. (2022b). For example, gender-biased word embeddings are known to embed the gender-neutral occupational words such as secretary, nurse, housekeeper, etc. such that they have high similarities with female pronouns than male pronouns Kaneko and Bollegala (2021). Therefore, a biased text encoder can retrieve documents that express stereotypically-biased opinions as supporting evidence for a query that does not explicitly mention any social biases. To evaluate this effect, we use three different retrieval methods in § 4.5.
Biases in the Generator:
An LLM acts as the generator in RAG, which generates a response considering both the query as well as the set of retrieved documents following a user-specified instruction. Even when the query and the retrieved documents are not biased, the social biases encoded in the LLM can still result in a biased response. To study this effect, we evaluate multiple generator LLMs, trained on different pre-train language data and parameter sizes in § 4.3.
@@
Following the QA-based social bias evaluation approach proposed by Parrish et al. (2022), we evaluate the social biases in a RAG system based on its ability to correctly answer questions without reflecting any unfair stereotypical biases. A test instance in a BBQ dataset contains a question (presented in a negated or a non-negated format), an ambiguous context (evaluates RAG behaviour in cases where there is insufficient evidence from the context to provide an answer) and a disambiguated context (provides information about which of the individuals mentioned in the ambiguous context is the correct answer). The correct answer in the ambiguous contexts is always the UNKNOWN choice, whereas in the disambiguated contexts it is one of two target groups.
Accuracy for the ambiguous contexts, $\textrm{Acc}_{\textrm{a}}$, is defined as the fraction of the ambiguous contexts predicted as UNKNOWN, while the accuracy for the disambiguated contexts, $\textrm{Acc}_{\textrm{d}}$, is defined as the fraction of the correct prediction of the disambiguous contexts for the specific target group. Accuracy does not indicate the directionality of the bias (i.e. stereotypically biased towards the advantaged group vs. anti-stereotypically biased towards the disadvantaged group). Advantaged groups refer to demographic groups that historically had greater access to resources, opportunities, power, or social privilege, whereas disadvantaged groups are those who have historically had discrimination, stereotypes, or unequal resource distributions.
To address this, Jin et al. (2024) proposed the Diff-Bias score as the difference of accuracies for the biased and counter-biased cases (see Appendix A for the definition). A zero Diff-Bias score indicates that the model under evaluation is not socially biased, while a positive or negative Diff-Bias score indicates social biases towards advantaged or disadvantaged groups, respectively. We use both Diff-Bias and Accuracy in our evaluations. However, due to the limited availability of space, all accuracy-based results are shown in Appendix C.
We provide the same instruction to all LLMs for BBQ evaluations. Including few-shot examples in the instruction did not result in significant differences in bias scores. Therefore, we used a zero-shot prompt for evaluations. Further details of the instructions are provided in Appendix D.
@@
We construct a comprehensive document collection to study the manifestation of various social biases in a RAG setting. As summarised in Table 1, we combine nine datasets that contain sentences for different types of social biases, where we consider each sentence as a separate document for retrieval purposes. The final collection contains 66,695 documents and is refereed to as the full-set henceforth. Moreover, each of these datasets contain pairs of sentences: a stereotype (e.g. women don’t know how to drive) and an anti-stereotype (e.g. men don’t know how to drive). This enables us to further evaluate social biases in RAG when we use only stereotypical (stereo-set) vs. anti-stereotypical (anti-set) sentences as the document collection.
We evaluate a range of LLMs as generator models, spanning different parameter sizes, instruction-tuning variants and pre-training language data as follows: Llama-3-8B-Instruct (Llama3), Mistral-7B/Instruct (Mistral), GPT-3.5-turbo (GPT-3.5), Llm-jp-3.1-Instruct 1.8B/7B/13B (Llm-jp), Qwen2.5-3B/7B/14B (Qwen) base and instruction-tuned versions. We use OpenAI API for GPT-3.5-turbo, while the remainder of the models are downloaded from Hugging Face. [^1] https://huggingface.co
For document retrieval, we consider three methods: (a) VectorIndex from LlamaIndex with 1536-dimensional OpenAI text-embedding-ada-002 embeddings, (b) BM25, a sparse retriever available in LlamaIndex, and (c) contriever, a contrastively pre-trained dense retrieval system Izacard et al. (2021) that uses the facebook/contriever retrieval model.
@@
StereoSet (Nadeem et al., 2021b) 1,744 - 5,894 482
Redditbias (Barikeri et al., 2021) 4,065 2,553 2,553 26,948
CrowSPairs (Névéol et al., 2022) 261 182 1,016 222
@@
Bias Type Setting GPT-3.5 Llama3-8B-Inst. Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst.
Gender w/o RAG 5.16 / -9.33 5.65 / 1.59 10.02 / -3.67 3.77 / -7.34 -2.38 / -2.38
stereo-set 14.53 / 7.14 14.68 / -0.4 24.01 / 0.5 13.99 / -2.68 4.61 / 2.68
full-set 11.31 / -0.1 6.80 / -3.97 15.43 / -2.08 0.55 / -4.66 -3.08 / -5.95
anti-set 4.51 / -3.97 0.74 / -6.85 10.17 / -10.12 -4.51 / -8.93 -8.43 / -12.7
Age w/o RAG 41.79 / 5.92 31.25 / 8.32 30.52 / 3.42 38.02 / 7.34 18.02 / 8.59
stereo-set 32.61 / 8.97 27.66 / 10.71 35.87 / 3.15 38.56 / 7.01 18.72 / 9.35
full-set 29.67 / 6.63 19.67 / 4.13 30.52 / 3.75 27.53 / 6.96 7.50 / 6.09
anti-set 17.83 / 6.30 8.97 / 2.77 20.11 / 3.53 6.96 / 6.79 2.69 / 3.26
Race w/o RAG 10.00 / 3.40 6.60 / 1.06 1.60 / 2.02 6.81 / 2.13 0.00 / -3.30
stereo-set 24.95 / 13.30 17.55 / 9.26 12.55 / 6.17 19.95 / 8.09 3.88 / 3.83
full-set 16.60 / 8.83 12.18 / 6.91 7.98 / 4.89 13.46 / 3.83 0.00 / 0.64
anti-set 6.49 / 5.43 7.23 / 4.15 4.73 / 6.11 4.36 / 2.23 -0.43 / -1.17
Religion w/o RAG 8.92 / 4.33 18.76 / 7.17 5.92 / 3.50 12.58 / 5.00 8.17 / 2.83
stereo-set 14.83 / 12.50 17.67 / 8.50 16.67 / 5.83 22.67 / 7.17 10.42 / 9.33
full-set 8.00 / 9.00 10.17 / 9.17 12.83 / 5.50 12.58 / 8.83 8.42 / 4.17
anti-set 2.83 / 5.17 11.92 / 8.83 6.50 / 3.67 8.00 / 5.00 7.75 / 2.00
Table 2: Diff-Bias scores for the ambiguous and disambiguated contexts (separated by ‘/’) for different bias types and models, with document collections of varying social bias levels used for retrieval. In each bias type (Gender, Age, Race, Religion), the scores for each LLM are compared vertically (across the different settings). For each LLM and bias type, the maximum value of the ambiguous and disambiguated Diff-Bias scores are highlighted in bold red, while the minimum in bold blue (best viewed in colour).
@@
Table 2 shows the Diff-Bias scores for the ambiguous and disambiguated contexts on the English BBQ dataset for gender, age, race and religion related social biases for four generator LLMs. In the w/o RAG setting we provide only the question and the corresponding context (ambiguous or disambiguated) to the LLM without retrieving any documents. This baseline shows the level of social biases in a generator LLM in the absence of RAG. On the other hand, full-set, stereo-set and anti-set methods use VectorIndex to retrieve the top-$10$ documents respectively from the full-set, stereo-set and anti-set document collections.
Overall, we see that full-set and stereo-set increase the social biases towards the advantaged group in each LLM compared to w/o RAG. In particular, we see that stereotypically biased documents (i.e stereo-set) result in the largest positive increases in social biases. On the other hand, anti-stereotypical documents (i.e. anti-set) often pushes the social biases in the opposite (towards the disadvantaged group) relative to w/o RAG. This result shows the high sensitivity of social biases in RAG to the external document collections. Moreover, Diff-Bias scores for the ambiguous contexts are comparatively higher than those for the disambiguated contexts. This indicates that, in the absence of informative contexts, LLMs tend to generate biased responses reflecting their internal social biases.
Among the four social bias types, we find that gender- and race-related biases, although relatively low in the baseline (w/o RAG) setting, are substantially amplified when the generator LLM retrieves documents from the stereo-set. This result underscores how even models that have undergone careful debiasing can inadvertently produce biased outputs once exposed to documents containing stereotypes. In contrast, biases pertaining to age and religion tend to be more pronounced in the original LLMs and exhibit only moderate increases under RAG.
Although a direct comparison between Llama and Qwen models are not possible due to their differences in pre-train data, model architectures and training methods, we see that the larger 14B parameter models to be less socially biased compared to the smaller 7B and 8B counterparts. This observation aligns with prior findings suggesting that larger LLMs often show reduced bias Zhou et al. (2023, 2024). Between the base Qwen-14B and the instruction tuned Qwen-14B-Instruct models, we see that the latter demonstrates lower Diff-Bias score for the ambiguous contexts. Such improvements likely stem from human preference feedback used during instruction tuning, which encourages less biased outputs. Unfortunately, as our results show, this safety alignment can be compromised once the instruction-tuned model is paired with a document collection that contains stereotypical content in a RAG pipeline.
@@
Llama3-8B-Inst. 5.65 / 1.59 14.68 / -0.40 6.80 / -3.97 0.74 / -6.85
Mistral 3.82 / 2.88 2.63 / 1.19 -2.53 / -3.37 -3.72 / -0.79
Mistral-Inst. -2.83 / 0.50 6.30 / 14.09 0.69 / 0.50 -10.47 / -0.40

### v2

@@ Abstract
Retrieval Augmented Generation (RAG) is a popular method for injecting up-to-date information into Large Language Model (LLM)-based Natural Language Generation (NLG) systems. While RAG can enhance factual accuracy, its effect on the social biases inherent in LLMs is not well understood. This paper systematically investigates how RAG modulates social biases across three languages (English, Japanese, and Chinese) and four categories (gender, race, age, and religion). By evaluating various generator LLMs on the BBQ benchmark, we analyse how document collections with controlled stereotypical content affect RAG outputs. We find that biases present in the retrieved documents are often significantly amplified in the generated texts, even when the base LLM itself has a low-level of intrinsic bias. These findings raise concerns about the social fairness of RAG systems, underscoring the urgent need for careful bias evaluation before real-world deployment.
@@
RAG Lewis et al. (2020); Edge et al. (2024) has become a cornerstone for enhancing LLMs, effectively mitigating hallucinations and incorporating up-to-date knowledge (Edge et al., 2024; Izacard and Grave, 2021). However, as LLMs are increasingly integrated into applications, a critical question remains: how does RAG interact with the social biases inherent in both pre-trained models and external knowledge sources?
@@
Social bias in NLP is defined in many, sometimes conflicting, ways (Blodgett et al., 2020). Following Parrish et al. (2022), we study the bias in Question Answering (QA), where a model’s discrete outputs manifest identifiable biases. Despite extensive evaluation of RAG systems for retrieval efficacy Wu et al. (2024); Laban et al. (2024); Yang et al. (2024) and factual accuracy Krishna et al. (2024); Soman and Roychowdhury (2024), their role in propagating social biases has been under explored. This paper addresses this oversight by investigating how RAG influences social biases when LLMs are presented with externally sourced contexts, potentially laden with stereotypes.
In the context of social biases, a stereotype is a widely held but oversimplified and fixed belief or assumption about the characteristics, attributes, or behaviours of members of a particular social group. Social groups can be further categorised into advantaged or disadvantaged. (Nangia et al., 2020) Advantaged groups refer to demographic groups that historically had greater access to resources, opportunities, power, or social privilege, whereas disadvantaged groups are those who have historically had discrimination, stereotypes, or unequal resource distributions. Anti-stereotypes are positive stereotypes held for the advantaged group.
We analyse the bias propagation using Bias Question Answering (BBQ) Parrish et al. (2022), a QA-structured benchmark that assesses social biases in LLMs, across gender, age, race and religion applying three retrieval methods. Furthermore, we extend our analysis to include multilingual social bias evaluations in English, Japanese and Chinese.
Our comprehensive analysis includes several key insights:
@@
    Across four types of social biases, retrieving documents labelled with stereotypes significantly amplifies bias in the final output, even for generator LLMs that initially exhibit low intrinsic biases.
@@
    Overall, social biases are less affected by the retrieval methods in RAG, while sparse retrieval tend to be more sensitive to social biases than the denser ones. Surprisingly, social biases are not monotonously increasing with the number of documents retrieved due to the decreasing relevance.
@@
    Bias amplification in RAG is not limited to English but is also a significant challenge in Chinese and Japanese languages. Even advanced RAG systems such as Self-RAG are not immune to this issue.
  4. 4.
    We evaluate potential solutions and find that while alignment tuning techniques such as Direct Preference Optimisation (DPO) offer a robust way to mitigate social biases. Social bias mitigation methods that use prompt-based in-contextual learning is less effective because of the increased context arising from the retrieved biased documents.
@@
Retrieval Augmented Generation
RAG Lewis et al. (2020); Edge et al. (2024) augments the inputs to LLMs with a set of relevant text passages/documents to mitigate missing knowledge to improving the accuracy of the responses. Modern RAG frameworks give the LLM a greater agency over the retrieval and generation process such as Self-RAG Asai et al. (2024) and CRAG Yan et al. (2024), which enable adaptive retrieval and self-correction, while RankRAG Yu et al. (2024) unifies context ranking with generation within a single model. This trend towards complexity is epitomised by architectures such as ModularRAG Gao et al. (2024), which involves multiple control components for retrieval and generation. Although powerful, the complex nature of these architectures makes it difficult to isolate the source of issues such as bias amplification. Consequently, in our study we use a standard, decomposable RAG pipeline as a deliberate methodological choice to enable a controlled, component-wise analysis of social bias propagation.
@@
Although social biases in LLMs have been studied extensively for various downstream applications, a RAG system can generate biased and harmful responses due to the social biases in the retrieved documents (Blodgett et al., 2020; Ni et al., 2025). Previous work measures gender-related biases in ranked lists and assesses the impact of different retrievers (Rekabsaz and Schedl, 2020). Kim et al. (2025) compared the effect of different fine-tuned encoders in RAG for gender-related biases. Hu et al. (2024) studied the fairness degradation in LLMs using datasets with different sub-categories of social biases and showed that even fully censored corpora can still induce biased responses. Wu et al. (2025) focused on fairness disparities in retrieval performance, analysing performance differences between protected and non-protected demographic groups.
However, these foundational studies are primarily limited to English and often utilize template-based contexts or a single collection of texts, which may not fully capture the complexity of real-world documents. Our work extends these works in several important directions: (i) decoupling the effects of the three core RAG components (document collection, retriever, and generator); (ii) employing controllable corpora with explicit stereotypical and anti-stereotypical polarities to precisely measure any bias amplifications; and (iii) extending the evaluation to a multilingual context and question sets (e.g. English, Chinese, and Japanese) across a richer set of human-annotated social biases.
Figure 2: Overview of our RAG social bias evaluation protocol. Given a collection of documents, encoded individually using an external encoder $f$, a vector index is created over the collection of the documents. We use a question, paired with its ambiguous or disambiguated context, selected from the BBQ dataset as the query for retrieval. We retrieve the top $k$ nearest neighbour documents to the query from the vector index, and provide them to the generator LLM, $g$, alongside with the question and the context.
@@
§ 3.1 RAG Framework
As illustrated in Figure 2, a standard RAG system consists of three core components: a Document Collection ($\mathcal{D}$) that serves as an external knowledge source, a Retriever ($f$) that indexes documents and fetches the top-$k$ relevant passages for a given query $q$, and a Generator ($g$), typically an LLM, that produces a response based on both the query and the retrieved documents. We systematically analyse how social biases can get amplified at each of those stages.
Biases in Documents:
The document collection $\mathcal{D}$ itself can contain stereotypical content. If there are many documents that express various levels of stereotypical social biases, then a subset of those documents can be retrieved even when the query does not explicitly mention any social biases and in return influence the generator to produce a biased response. We evaluate this by using document collections with controlled levels of stereotypical and anti-stereotypical content on four types of social biases (i.e. gender, age, race, religion) (§ 4.2) in the document collection using three benchmark datasets covering English, Japanese Yanaka et al. (2024) and Chinese Huang and Xiong (2024) languages in § 4.5.
Biases in the Retriever:
The text encoder used by the retriever $f$ can encode unfair social biases, causing it to preferentially retrieve documents that confirm stereotypes, even for neutral queries Bolukbasi et al. (2016); Kim et al. (2025). We investigate this effect by comparing different retrieval methods in § $4.$
Biases in the Generator:
The generator $g$ may have its own biases learned during pre-training. These can manifest in the final output, even if the retrieved documents are unbiased. We study this effect by evaluating multiple LLMs of varying sizes and training data in § 4.3.
@@
We adopt the Bias Benchmark for Question Answering (BBQ) dataset Parrish et al. (2022) to quantify social biases. BBQ provides a controlled setting to distinguish between a model’s reliance on stereotypes versus its ability to follow explicit evidence. Each BBQ instance contains a question and two types of contexts:
Ambiguous Contexts:
There is insufficient information to answer the question, making “UNKNOWN” the only unbiased response. For example, in Figure 2, given “We couldn’t start the board meeting at 9 am today because a man and a woman were late”, we must answer “Who was the secretary?”. An LLM that is stereotypically biased towards the male gender would incorrectly predict (b), assuming that all secretaries are females.
Disambiguated Contexts:
@@
An LLM which always selects (c) and refrain from answering any socially sensitive questions would appear to be unbiased, when in fact it might not be. Therefore, the disambiguated contexts explicitly provide clear evidence to determine the correct answer, testing the model’s ability to prioritise given facts over internal biases. In the previous case, provided the disambiguation context in BBQ, “We had to wait for the woman because they were running the meeting, and the man was responsible for taking all the notes”, the correct answer to this question would be (a). This QA format allows us to evaluate bias through discrete choices, avoiding the complexities of analysing open-ended generation.
@@
Following the QA-based social bias evaluation approach proposed by Parrish et al. (2022), we evaluate the social biases in a modular RAG system based on its ability to correctly answer questions without reflecting any unfair stereotypical biases.
Accuracy
measures the model’s performance. For ambiguous contexts ($\textrm{Acc}_{\textrm{a}}$), it is the fraction of “UNKNOWN” answers. For disambiguated contexts ($\textrm{Acc}_{\textrm{d}}$), it is the fraction of correct answers based on the provided evidence. However, accuracy does not indicate the directionality of the bias (i.e. stereotypically biased towards the advantaged group vs. anti-stereotypically biased towards the disadvantaged group).
Diff-Bias (Jin et al., 2024)
addresses this by measuring the difference between stereotypical and anti-stereotypical responses(see Appendix A for the definition). A score of zero indicates no bias, while positive or negative scores indicate bias towards advantaged or disadvantaged groups, respectively. We report Diff-Bias for both ambiguous and disambiguated contexts. Due to space constraints, full accuracy results are presented in Appendix F.
@@
We construct a comprehensive document collection to study the manifestation of various social biases in a RAG setting. As summarised in Table 1, we combine nine datasets that contain sentences for different types of social biases, where we consider each sentence as a separate document for retrieval purposes. The final collection contains 64,142 documents and is refereed to as the full-set henceforth. Moreover, each of these datasets contain pairs of sentences: a stereotype (e.g. women don’t know how to drive) and an anti-stereotype (e.g. men don’t know how to drive). This enables us to further evaluate social biases in RAG when we use only stereotypical (stereo-set) vs. anti-stereotypical (anti-set) sentences as the document collection.
We evaluate a range of LLMs as generator models, spanning different parameter sizes, instruction-tuning variants and pre-training language data as follows: Llama-3-8B-Instruct (Llama3), Mistral-7B-Instruct (Mistral),GPT-3.5-turbo (GPT-3.5), Qwen2.5-3B/7B/14B/72B (Qwen) base and instruction-tuned versions. We use OpenAI API for GPT-3.5-turbo and o4-mini, while the remainder of the models are downloaded from Hugging Face. [^1] https://huggingface.co For document retrieval, we consider three methods: (a) VectorIndex from LlamaIndex with 1536-dimensional OpenAI text-embedding-ada-002 embeddings, (b) BM25, a sparse retriever available in LlamaIndex, and (c) Contriever, a contrastively pre-trained dense retrieval system Izacard et al. (2021) that uses the facebook/contriever retrieval model.
@@
StereoSet (Nadeem et al., 2021) 1,744 - 5,894 482
Redditbias (Barikeri et al., 2021) 4,065 - 2,553 26,948
CrowSPairs (Nangia et al., 2020) 261 182 1,016 222
@@
GenderAgeRaceReligion01010202030304040Diff-BiasAmbiguous Contextsw/o RAGstereo-setfull-setanti-set
GenderAgeRaceReligion9-96-63-30336699Disambiguated Contexts
Figure 3: Diff-Bias under different retrieval sets (averaged over 6 LLMs). Bars show the mean Diff-Bias across GPT-3.5, Llama3-8B-Inst., Qwen-7B-Inst., Qwen-14B base and Inst. and Qwen-72B-Inst. for each bias type. Gray=w/o RAG, red=stereo-set, green=full-set, blue=anti-set. Error bars representing Confidential Intervals (CIs) are omitted for visual clarity.(see Table 17 for full CIs for each model and bias category)
@@
In this section, we focus on how the polarity of the document collection influences bias. Figure 3 presents the average Diff-Bias scores for the ambiguous and disambiguated contexts on the English BBQ dataset for gender, age, race and religion related social biases across six generator LLMs.
Document Polarity is the main Factor.
The bias polarity of the retrieved documents is the main factor driving the final output bias, as is shown in Figure 3. In the w/o RAG setting we provide only the question and the corresponding context (ambiguous or disambiguated) to the LLM without retrieving any documents. On the other hand, full-set, stereo-set and anti-set methods use VectorIndex to retrieve the top-$10$ documents respectively from the corresponding document collections.
Compared to the w/o RAG baseline (gray bars), the stereo-set (red bars) amplifies bias towards the advantaged group across all four bias categories. Conversely, the anti-set (blue bars) reduces the Diff-Bias score, often pushing the biases in the opposite direction, which indicates that the bias is not merely mitigated but reversed towards the disadvantaged group. This shows the high sensitivity of biases in RAG to the external documents.
The Critical Role of Ambiguity.
We also find that Diff-Bias for the ambiguous contexts (Figure 3, left) are higher than that for the disambiguated contexts (Figure 3, right). The effect of document polarity is more pronounced in ambiguous scenarios. When the model is provided with informative evidence in disambiguated contexts, the influence of the document collection’s bias is highly reduced. This highlights that RAG poses the fairness risk in information-lacking, ambiguous situations, relying on the biases in the retrieved documents.
Varying Amplification across Bias Types.
Among the four social bias types, we find that gender- and race-related biases, although relatively low in the baseline (w/o RAG), are substantially amplified when the generator retrieves documents from the stereo-set. The pattern of non‐overlapping 95% CIs (see Table 17) further confirms that stereo‐set consistently produces the largest increases in Diff‐Bias on ambiguous questions. In contrast, age-related biases are notably higher in the baseline models themselves, a finding consistent with prior work suggesting that LLMs often encode negative stereotypes about older individuals (Shin et al., 2024). While RAG still amplifies this bias, the relative increase is more moderate compared to the baseline.
The Risk of Intersectional Bias:
Prior work has shown that social bias categories are correlated and often results in intersectional biases across categories Tan and Celis (2019); Lalor et al. (2022); Ma et al. (2023). This has important implications for RAG where the queries and documents cover different social bias categories, even when the targeted bias category has been filtered or is absent from the retrieved documents. We found that such intersectional biases are also consistently amplified during RAG (see Appendix D), which raises serious concerns.
@@
Llama3-8B-Inst. 5.65 / 1.59 14.68* / -0.40 6.80 / -3.97 0.74 / -6.85
Mistral-Inst. -2.83 / 0.50 6.30* / 14.09* 0.69 / 0.50 -10.47 / -0.40